OpenStudy (anonymous):

guys i have an important quiz tomorrow and i need help asap ... The position of a particle in meters is given by, x(t)=16t-3.0t^3 where the time t is in seconds. At what time is the particle momentarily at rest ?

OpenStudy (anonymous):

the particle is at rest then it means velocity is zero, so v=0=> dx(t)/dt=0 x(t)=16t-3.0t^3 dx(t)/dt=16-9t^2=0 t^2=16/9=>t=4/3sec

OpenStudy (anonymous):

check the answer urself...all the best